OpenStudy (anonymous):

If you want your retirement account to have a balance of $200000 when you retire in 50 years, how much should you invest now at a rate of 8% interest compounded continuously?

OpenStudy (amistre64):

A = Pe^(rt) solve for P and input the given values from the question

OpenStudy (amistre64):

or input then solve ... either way

OpenStudy (anonymous):

I did and got it to equal 3.05 as the initial value.. which the book says is wrong

OpenStudy (amistre64):

can you tell me what you inputed for the values of A, r, t ?

OpenStudy (anonymous):

i made a=200,000 r=.08 t=50

OpenStudy (amistre64):

those are good inputs .... must be in the solving for P process \[A=Pe^{rt}\] \[A/e^{rt}=P\] \[200000/e^{.08*50}=P\] http://www.wolframalpha.com/input/?i=200000%2Fe%5E%28.08*50%29

OpenStudy (anonymous):

Thank you!